OpenStudy (anonymous):

an angle who supplement is 20 degrees more than twice its complement

OpenStudy (anonymous):

180-x=20+2(90-x) 180-x=20+180-2x 180-x=200-2x -x + 2x =200-180 x = 20 angle is 20 degree If I deserve medal then give me one

OpenStudy (anonymous):

?

OpenStudy (anonymous):

first time here, how do I give out medals

OpenStudy (anonymous):

next to my name click give medal

OpenStudy (anonymous):

well i must be having a moment because I don't see anything that say give medal

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

thank, how u do it?

OpenStudy (anonymous):

I just click where it said good answer

OpenStudy (anonymous):

ok thank